if secx+tanx=5, then find the value of (tanx+secx-1)/(tanx-secx+1) 
 

Grade:10

1 Answers

Hari Sundar Pyda
6 years ago
secx+tanx=5(tanx+secx-1)/(tanx-secx+1)= (tanx+5-tanx-1) / (tanx+5-tanx+1)= 4 / 6= 2 / 3.Hope it helps!!